# Evaluation of Google Translate for Mandarin Chinese translation using sentiment and semantic analysis

### v1

@@ Abstract
Machine translation using large language models (LLMs) is having a significant global impact, making communication easier. Mandarin Chinese is the official language used for communication by the government, education institutes, and media in China. In this study, we provide an automated assessment of machine translation models with human experts using sentiment and semantic analysis. In order to demonstrate our framework, we select classic early twentieth-century novel ’The True Story of Ah Q’ with selected Mandarin Chinese to English translations. We also us Google Translate to generate the given text into English and then conduct a chapter-wise sentiment analysis and semantic analysis to compare the extracted sentiments across the different translations. We utilise LLMs for semantic and sentiment analysis. Our results indicate that the precision of Google Translate differs both in terms of semantic and sentiment analysis when compared to human expert translations. We find that Google Translate is unable to translate some of the specific words or phrases in Chinese, such as Chinese traditional allusions. The mistranslations have to its lack of contextual significance and historical knowledge of China. Thus, this framework brought us some new insights about machine translation for Chinese Mandarin. The future work can explore other languages or types of texts with this framework.
@@ 4.2 Sentiment analysis
Additionally, we explore the core sentiment of each chapter leading towards either positive or negative sentiments, thus we visualise the sentiment polarity score of each chapter in Figure LABEL:fig:polarity_hsien. We categorise the value of each sentiment as either 1 or -1, depending on whether it is a positive sentiment or a negative sentiment. We assign sentiments ”optimistic”, ”thankful” and ”empathetic” with a positive sentiment score of + 1. In contrast, we assign sentiments such as ”pessimistic”, ”anxious”, ”sad”, ”annoyed” and ”denial” with a negative score of - 1. Then, we assign ”Joking” with a score of 0 since it could be treated as either a positive or negative sentiment. In Figure LABEL:fig:polarity_hsien, we note that the polarity score in all chapters is negative, and shows an overall declining trend. The plot initially shows a downward trend until chapter three, and there is a slight increase from chapter three to chapter five, followed by another downward trend until the end. The slight increase from chapter three to chapter five is because although Ah Q experienced three failures in total across three chapters, he also had some successes. Those successes in the storyline seem to influence the positive sentiment.
@@ 4.3 Semantic analysis
Chapter 2 0.645 0.591 0.764
Chapter 3 0.633 0.578 0.725
Chapter 4 0.711 0.645 0.724
Chapter 5 0.681 0.634 0.764
Chapter 6 0.698 0.540 0.587
Chapter 7 0.680 0.617 0.691
Chapter 8 0.700 0.619 0.728
Chapter 9 0.736 0.681 0.796

### v2

@@ Abstract
Machine translation using large language models (LLMs) is having a significant global impact, making communication easier. Mandarin Chinese is the official language used for communication by the government and media in China. In this study, we provide an automated assessment of translation quality of Google Translate with human experts using sentiment and semantic analysis. In order to demonstrate our framework, we select the classic early twentieth-century novel ’The True Story of Ah Q’ with selected Mandarin Chinese to English translations. We use Google Translate to translate the given text into English and then conduct a chapter-wise sentiment analysis and semantic analysis to compare the extracted sentiments across the different translations. Our results indicate that the precision of Google Translate differs both in terms of semantic and sentiment analysis when compared to human expert translations. We find that Google Translate is unable to translate some of the specific words or phrases in Chinese, such as Chinese traditional allusions. The mistranslations may be due to lack of contextual significance and historical knowledge of China.
@@ 4.2 Sentiment analysis
In Figure 5 and Figure 6, we observe that [foreign, devil] and [successful, provincial] is leading permutation in joking and annoyed bi-grams for Xianyi and Gladys, while [foreign, devil] and [fake, foreign] is leading words in joking and annoyed bi-grams for Julia. Moreover, by evaluating joking and annoyed tri-grams for all translations, it is interesting to note that [imitation, foreign, devil] is leading words in both joking and annoyed sentiments for Xianyi and Gladys, and [fake, foreign, devil] is leading words in both joking and annoyed sentiments for Julia. It further illustrates that the expression in terms of joking and annoyed by two human experts translation are generally similar. Thus, it gives a message, that the cumulative counts of ’Annoyed’ and ’Joking’ sentiments in the two translations shows no huge differences in Figure 7.
Additionally, we explore the core sentiment of each chapter leading towards either positive or negative sentiments, thus we visualise the sentiment polarity score of each chapter in Figure LABEL:fig:polarity_hsien. We categorise the value of each sentiment as either 1 or -1, depending on whether it is a positive sentiment or a negative sentiment. We assign sentiments ”optimistic” with a positive score of + 2, ”thankful” with a score of + 3 and ”joking” with a score of + 1 as they are positive sentiments. In contrast, we assign sentiments ”pessimistic” with a score of -4, ”anxious” with a score of -2, ”sad” with a score of -3, ”annoyed” with a score of -1 and ”denial” with a score of -5. Then, we assign ”Empathetic” with a score of 0 since it could be treated as either positive or negative sentiments. By observing Figure 8, it is interesting to note that the mean of polarity score across all chapters are mostly negative (except Google Translate translation). The trend of two human experts translation represent similar features, which both start with an increasing trend, remain relatively steady trend in middle chapters, then show an upward trend from chapter 6 to chapter 7 and end with a downward trend. This further demonstrates that the two human experts translations have a similar overall understanding for the original text in sentiment perspective. However, the mean of polarity for Google Translate translation has different trend compared with two human experts translation. For example, the translation by Google Translate shows a obvious downward trend from chapter 2 to chapter 3, whereas the feature is not observed from two human translations. The polarity score of Google Translate translation deviates from the two human experts translations. Figure 9 is used a textual data processing library in Python known as TextBlob. The range of polarity score in TextBlob is given from -1 to 1. By comparing mean of polarity score by first method and mean of polarity score by TextBlob, the range of the mean polarity score by TextBlob is higher than the mean polarity score by the first method. Moreover, in Figure 9, the mean of polarity scores across three translations are different. We notice that both Google Translate translation and Julia translation has a lowest point in chapter 4 whereas the translation by Xianyi and Glady does not exhibit this feature. It is interesting to note that the translations by two human experts show a downward trend from chapter 8 to chapter 9 in both Figure 8 and Figure 9. This reflects that Ah Q involved in the revolution and ultimately becomes a victim for the revolution and being executed in chapter 9, leading to a tragic end.
@@ 4.3 Semantic analysis
Chapter 2 0.645(0.148) 0.591(0.139) 0.764(0.124)
Chapter 3 0.633(0.139) 0.578(0.159) 0.725(0.124)
Chapter 4 0.711(0.139) 0.645(0.155) 0.724(0.163)
Chapter 5 0.681(0.117) 0.634(0.161) 0.764(0.140)
Chapter 6 0.698(0.106) 0.540(0.192) 0.587(0.215)
Chapter 7 0.680(0.174) 0.617(0.180) 0.691(0.161)
Chapter 8 0.700(0.144) 0.619(0.166) 0.728(0.133)
Chapter 9 0.736(0.128) 0.681(0.130) 0.796(0.117)